OpenStudy (meowmeowman):

a chicken starts running eastwar from a point 30 meters due west of a telephone pole. The chicken runs at 4.1 meters per second. At the same instant, a gopher starts from a point 40 meters north of the pole, heading south at 3.8 meters per second. After how much time will the chicken and gopher first be 20 meters apart?

OpenStudy (anonymous):

Why is the chicken running?

OpenStudy (anonymous):

Draw picture.

OpenStudy (meowmeowman):

I did

OpenStudy (meowmeowman):

stil dont get it

OpenStudy (anonymous):

The chicken's position is given by: \[ f(t)=4.1t \]The gopher's position is given by: \[ g(t)=40-3.8t \]

OpenStudy (anonymous):

The distance between them is \[ d(t) = \sqrt{[f(t)]^2+[g(t)]^2} \]

OpenStudy (meowmeowman):

the chicken is 30 meters from the pole

OpenStudy (anonymous):

You need to solve for \(t\) where: \[ 20=\sqrt{(30+4.1t)^2+(40-2.8t)^2} \]

OpenStudy (anonymous):

Oh wait...

OpenStudy (anonymous):

You need to solve for \(t\) where: \[ 20=\sqrt{(30-4.1t)^2+(40-2.8t)^2} \]

OpenStudy (anonymous):

You need to find the lowest \(t\) where \(t>0\)

OpenStudy (anonymous):

I get \(t\approx 7.15\)

OpenStudy (meowmeowman):

i know the answer, but dont know how to do it. And by the way, it is 5.6 seconds

OpenStudy (anonymous):

Now I'm getting \(t=5.6\) using: \[ 20=\sqrt{(30-4.1t)^2+(40-3.8t)^2} \]

OpenStudy (anonymous):

\(\color{blue}{\text{Originally Posted by}}\) @wio The distance between them is \[ d(t) = \sqrt{[f(t)]^2+[g(t)]^2} \] \(\color{blue}{\text{End of Quote}}\) This is the essential concept to use.

OpenStudy (anonymous):

You just need to model the position of each creature.

OpenStudy (meowmeowman):

so you endded up getting 5.6 seconds with the formula// ?

OpenStudy (anonymous):

Yes.

OpenStudy (meowmeowman):

thank you very much! I raelly needed it

OpenStudy (anonymous):

Okay, but I don't know why chicken is running.

OpenStudy (meowmeowman):

the world may never know